OpenStudy (sphott51):

solve 15^2x = 36 log

OpenStudy (nincompoop):

obtain the Ln of both sides

OpenStudy (sphott51):

log(15^2x) = log(36) ?

OpenStudy (nincompoop):

logarithmic base 15 of both sides

OpenStudy (sphott51):

2xlog (15) =log (36) ?

OpenStudy (sphott51):